when a hydrogen sample in ground state is bombarded then what potential is required to accelerate electron so that first paschen line is emmited

Grade:12th pass

1 Answers

Arun
6 years ago
The energy of electronic levels in hydrogen atom is determined by the equation:
En=-13.6/n2
, where n – the level number.
The first Paschen line corresponds the transition n=4 to n=3. Thus, the electron should be delivered to 
n=4 level. The potential required to do this equals:
E = E4-E1 = 13.6(1 – 1/16) eV = 12.75 eV